solvinng system of equation using three variables
6x-y+z=2
4x+2y-3z=18
x-3y+2z=-5

Answer by mananth(16949)

6x-y+z = 2-------------- 1
4x+2y-3 z=18--------------2
x-3y+2z=-5--------------3
consider equation 1 &2 Eliminate y
Multiply 1 by 2
Multiply 2 by 1
we get
12x-2y +2z=4
4x+2y-3z=18
Add the two
16x-1z=22-------------4
consider equation 2 & 3
Multiply 2 by 3
Multiply 3 by 2
we get
12x+6y-9z=54
2x-6y+4z=-10
Add the two
14x-5z=44 -------------5 5
Consider (4) & (5)
Multiply 4 by -14
Multiply (5) by 16
we get
-224x+14z=-308
224x-80z=704
Add the two
-66z=396
/-66
z=-6

Plug the value of z in 5
14x+30 =44
14x=14
x=1
plug value of x & z in 1
6-1y-6= 2
-y=-6+6+2
-y=2
y=-2